OpenStudy (anonymous):

when p(x)=x^2 + ax -b is divided by x-1, the remainder is 6. when p(x) is divided by x+2. the remainder is 3. find the values of a and b.

terenzreignz (terenzreignz):

Combines polynomials with systems of linear equations. Like I told Faye, the way to get the remainder when f(x) is divided by (x-c) is to simply plug in c into the function. IE Remainder of p(x) when it's divided by x-1... is just p(1) So... what is p(1)

OpenStudy (anonymous):

x-c

terenzreignz (terenzreignz):

wth?

OpenStudy (anonymous):

p(x)=x^2+ax-b

terenzreignz (terenzreignz):

mhmm so p(1) = ?

OpenStudy (anonymous):

i was just asking what x-c is.

terenzreignz (terenzreignz):

Just an example.

OpenStudy (anonymous):

oh. xDD

terenzreignz (terenzreignz):

Or, a formula. In this case, x-c is x-1

terenzreignz (terenzreignz):

So, c=1 now just plug in 1 into the function p(x) = x² + ax - b

OpenStudy (anonymous):

okaay,, p(1)=(1)^2+a(1)-b

OpenStudy (anonymous):

after that??

terenzreignz (terenzreignz):

Simplify -.- 1² + a(1) - b = 1 + a - b Okay, and this is equal to -6, as per the given. -6 = a - b + 1

OpenStudy (anonymous):

sorry,, typo,, its not -6 its only 6

terenzreignz (terenzreignz):

hmm... 6 = a - b + 1 Understood?

terenzreignz (terenzreignz):

Or, simplify further, we get a - b = 5

terenzreignz (terenzreignz):

... stay with me here...

OpenStudy (anonymous):

why 6 = a - b + 1??? where did you get +1??

terenzreignz (terenzreignz):

Focus... since the remainder of p(x) when it's divided by x-1 is 6, it can only mean that p(1) = 6 p(1) = 1² + a(1) - b = 1 + a - b etc

OpenStudy (anonymous):

ahah, xDD sorry,, didn't see thatxDD

OpenStudy (anonymous):

after that??

terenzreignz (terenzreignz):

You know a-b = 5 That's not enough to know the value of a and b, you need another equation. Now that we mention it... we have another fact... that when it's divided by x + 2, the remainder is 3. Can only mean that p(-2) = 3 Solve this part.

terenzreignz (terenzreignz):

Hey... p(-2)... what is it?

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

sorry,, we just started doing that today, xDD

terenzreignz (terenzreignz):

No better time than now. It's just plugging in x = -2... how hard can it be? :P p(-2) = ?

OpenStudy (anonymous):

p(-2)=(-2)^2+a(-2)-b

terenzreignz (terenzreignz):

simplify...

OpenStudy (anonymous):

4-2a-b

terenzreignz (terenzreignz):

And this is equal to...?

OpenStudy (anonymous):

2a-b=-4

terenzreignz (terenzreignz):

Excuse me...?

OpenStudy (anonymous):

my eyes hurts,.. T_T

terenzreignz (terenzreignz):

-.- my eyes *hurt* and anyway... you're losing focus...

OpenStudy (anonymous):

-2a-b=4

terenzreignz (terenzreignz):

Don't you just hate it when the other kid gets the right answer and then 'corrects' it with an incorrect one -.-

terenzreignz (terenzreignz):

Okay, no more guessing... show your work :P

OpenStudy (anonymous):

2a+b=1

terenzreignz (terenzreignz):

Show your work... let me know you're not guessing...

OpenStudy (anonymous):

(-2)^2+a(-2)-b=3 4-2a-b=3 -2a-b=3-4 -2a-b=-1 -2a-b=-1 divide by -1 2a+b=1

terenzreignz (terenzreignz):

Much better. Now you have both a - b = 5 and 2a + b = 1 I can trust you can solve this system with ease now?

OpenStudy (anonymous):

2a+b-1=0

terenzreignz (terenzreignz):

We certainly have a lot to work on -.- FIRST EQUATION a - b = 5 adding b to both sides gives a = b + 5 Now substitute b+5 for the a in the SECOND equation: 2(b+5) + b = 1 SOLVE FOR b. Get to it...

OpenStudy (anonymous):

a=2 b=-3, correct ba toh??

terenzreignz (terenzreignz):

Nice going, *peanut-butter boy* :P

OpenStudy (anonymous):

da eff???? how da eff,,

terenzreignz (terenzreignz):

pff... Google Derek Bocar... first result... You take so long, I have to find ways to ... kill time... and guess what came up? XD

OpenStudy (anonymous):

yeah,, that was really long time agoo, T_T

OpenStudy (anonymous):

most humiliating scandal

terenzreignz (terenzreignz):

mhmm... last bit just killed my appetite :D

terenzreignz (terenzreignz):

But enough about that. Job well done... you certainly picked this up with more speed than Faye...

OpenStudy (anonymous):

i'm really sleeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeepyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy!!!!!!

terenzreignz (terenzreignz):

Join the club. Why don't you listen to yourself playing? >:)

terenzreignz (terenzreignz):

LOL Just kidding Peace ^_^

OpenStudy (anonymous):

NOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

i'm gonna hack their youtube account!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

*his

terenzreignz (terenzreignz):

Come to think of it... just how good are you with that fiddle?

OpenStudy (anonymous):

ahm,, xDDD average?? xDDD not really an expert,

terenzreignz (terenzreignz):

figures... that's no use, I need to hear it. Any excerpts?

OpenStudy (anonymous):

i don't really play classical music, so long,,,, xDDD

OpenStudy (anonymous):

nope,, i don't really show my skills too much, xDD

terenzreignz (terenzreignz):

Le gasp... okay, interest over XD time to play some notes on a keyboard... ...the one that looks like a monster's smile and not the one with QWERTY written on it...

terenzreignz (terenzreignz):

<plays some random stuff>

terenzreignz (terenzreignz):

@derek11111 Listening to myself play is making me even sleepier... you should follow suit. Good night, -peanut butter boy- XD -------------------------------------- Terence out ^_^